# Covariance spectrum of MAXI J1820+070: On the nature of the Comptonizing flow

### v1

@@ 5 Acknowledgements
We thank F.M. Vincentelli, Yi Long, and Zhen Yan for their valuable discussion on spectral analysis. B.Y. is supported by Natural Science Foundation of China (NSFC) grants 12322307, 12361131579, and 12273026; by Xiaomi Foundation / Xiaomi Young Talents Program. AJT acknowledges that this research was undertaken thanks to funding from the Canada Research Chairs Program and the support of the Natural Sciences and Engineering Research Council of Canada (NSERC; funding reference number RGPIN-2024-04458). AAZ acknowledges support from the Polish National Science Center grants 2019/35/B/ST9/03944 and 2023/48/Q/ST9/00138.
@@
As mentioned in Section 3, we followed the procedure outlined in Uttley2014 to derive the coherence and covariance.
The Fourier cross-spectrum between two light curves x(t) and y(t) with DFTs $\rm X_{n}$ and $\rm Y_{n}$ is defined to be:
@@
For a given frequency bin $\nu_{j}$, the averaged cross-spectrum should be derived as follow:
@@
$\gamma^{2}(\nu)=\frac{|\bar{C}_{XY}(\nu)|^{2}-n^{2}}{(\bar{P}_{X}(\nu)-P_{X,\text{noise}})(\bar{P}_{Y}(\nu)-P_{Y,\text{noise}})},$ (3)
where $\bar{C}_{XY}(\nu)$ represents the average cross-spectrum within the specified frequency range. The term $\Delta\nu$ indicates the width of this frequency range. All power spectra and cross-spectra are normalized using Miyamoto normalization, which is adjusted to the fractional RMS according to Miyamoto1991.
@@
The error on the intrinsic coherence can be estimated as [Vaughan_1997, their equation 8]
$\Delta\gamma^{2}(\nu)=\gamma^{2}(\nu)\times\sqrt{[\frac{2n^{4}KM}{\left(|\bar{C}_{XY}(\nu)|^{2}-n^{2}\right)^{2}}+\frac{P_{X,\text{ noise }}^{4}}{\bar{P}_{X}\left(\nu\right)^{4}}+\frac{P_{Y,\text{ noise }}^{4}}{\bar{P}_{Y}\left(\nu\right)^{4}}+\frac{KM\delta\gamma^{2}}{\gamma^{4}}]/KM},$ (4)
@@
Following the approach outlined in Uttley2014, we derive covariance spectra over both long and short timescales using light curves from 50 specific energy channels, along with a reference band. For this study, we have chosen the 2-10 keV energy range as the reference band. For each energy channel of interest, we first construct the power spectra $P_{X}(\nu)$ and $P_{Y}(\nu)$ from the light curves of the energy channel and the reference band, which we denote as $x(t)$ and $y(t)$, respectively. It is important to note that the power spectra discussed here are not noise-subtracted. Using these two light curves, we also compute the cross-spectrum $C_{XY}(\nu)$. To generate the power and cross-spectra, we utilize the Python-based package Stingray. All power spectra and cross-spectra are normalized using Miyamoto normalization, which is based on the fractional RMS (Miyamoto1991). After deriving power spectra and the cross-spectrum, the value of the covariance spectrum in each energy channel of interest is given by:
$C_{v}\left(\nu\right)=\langle x\rangle\sqrt{\frac{\Delta\nu\left(\left|\bar{C}_{XY}\left(\nu\right)\right|^{2}-n^{2}\right)}{\bar{P}_{Y}\left(\nu\right)-P_{Y,\text{ noise }}}},$ (5)
where the first term $\langle x\rangle$ is the average count in the energy channel of interest. The second term is dimensionless; we refer to it as fractional covariance.
Note that the covariance spectrum is closely related to coherence. The equation can be expressed as follows:
$Cv\left(\nu\right)=\langle x\rangle\sqrt{\gamma^{2}\left(\nu\right)}\sqrt{\left(\bar{P}_{X}\left(\nu\right)-P_{X,\text{noise}}\right)\Delta\nu},$ (6)
where $\gamma^{2}$ represents the coherence between two light curves within the specified frequency range. Since the power spectrum utilizes fractional RMS-squared normalization, the term $\sigma_{\rm X}(\nu)\equiv\sqrt{\left(\bar{P}_{X}\left(\nu\right)-P_{X,\text{noise}}\right)\Delta\nu}$ indicates the fractional RMS of the energy channel within that frequency range. It is also essential to apply the background correction to this term, as given by the formula $RMS=\sqrt{P}(S+B)/S$ (Bu2015). Additionally, the error on the covariance could be derived via error propagation:
$\Delta Cv\left(\nu\right)=\sqrt{\left(\sqrt{\gamma^{2}(\nu)}\cdot\sigma_{\rm X}(\nu)\cdot\Delta{\rm\langle x\rangle}\right)^{2}+\left(\dfrac{\sigma_{\rm X}(\nu)\cdot\langle x\rangle\cdot\Delta\gamma^{2}(\nu)}{2\sqrt{\gamma^{2}(\nu)}}\right)^{2}+\left(\sqrt{\gamma^{2}(\nu)}\cdot\langle x\rangle\cdot\Delta\sigma_{\rm X}(\nu)\right)^{2}},$ (7)
where $\Delta\sigma_{\rm X}(\nu)$ is the error on the fractional rms-spectrum.

### v2

@@ V Acknowledgements
We thank F.M. Vincentelli, Yi Long, and Zhen Yan for their valuable discussion on spectral analysis. B.Y. is supported by Natural Science Foundation of China (NSFC) grants 12322307, 12361131579, and 12273026; by Xiaomi Foundation / Xiaomi Young Talents Program. The data analysis in this paper have been done on the supercomputing system in the Supercomputing Center of Wuhan University. AJT acknowledges that this research was undertaken thanks to funding from the Canada Research Chairs Program and the support of the Natural Sciences and Engineering Research Council of Canada (NSERC; funding reference number RGPIN-2024-04458). AAZ acknowledges support from the Polish National Science Center grants 2019/35/B/ST9/03944 and 2023/48/Q/ST9/00138.
@@
As mentioned in Section III, we followed the procedure outlined in Uttley et al. (2014) to derive the coherence and covariance.
The Fourier cross-spectrum between two light curves x(t) and y(t) with discrete Fourier transforms (DFTs) $\rm X_{n}$ and $\rm Y_{n}$ is defined to be:
@@
In a given frequency bin $\nu_{j}$, the averaged cross-spectrum over M segments and K frequencies per segment should be derived as follows:
@@
where $\bar{C}_{XY}\left(v_{j}\right)$ is the estimate of the cross-spectrum from the average of the periodogram in the bin $\nu_{j}$ and $C_{XY,n,m}$ is the value of a single sample of the periodogram measured from the $m$th segment with a frequency $f_{n}$ that is contained within the frequency bin $\nu_{j}$.
@@
$\gamma^{2}(\nu_{j})=\frac{|\bar{C}_{XY}(\nu_{j})|^{2}-n^{2}}{(\bar{P}_{X}(\nu_{j})-P_{X,\text{noise}})(\bar{P}_{Y}(\nu_{j})-P_{Y,\text{noise}})},$ (4)
where $\bar{C}_{XY}(\nu_{j})$ represents the average cross-spectrum within the specified frequency range. All power spectra and cross-spectra are normalized using Miyamoto normalization, which is adjusted to the fractional RMS according to Miyamoto et al. (1991).
@@
The error on the intrinsic coherence can be estimated as [Vaughan and Nowak (1997), their equation 8]
$\Delta\gamma^{2}(\nu_{j})=\gamma^{2}(\nu_{j})\times\sqrt{[\frac{2n^{4}KM}{\left(|\bar{C}_{XY}(\nu_{j})|^{2}-n^{2}\right)^{2}}+\frac{P_{X,\text{ noise }}^{4}}{\bar{P}_{X}\left(\nu_{j}\right)^{4}}+\frac{P_{Y,\text{ noise }}^{4}}{\bar{P}_{Y}\left(\nu_{j}\right)^{4}}+\frac{KM\delta\gamma^{2}}{\gamma^{4}}]/KM},$ (5)
@@
Following the approach outlined in Uttley et al. (2014), we derive covariance spectra over both long and short timescales using light curves from 50 specific energy channels, along with a reference band. For this study, we have chosen the 2-10 keV energy range as the reference band. For each energy channel of interest, we first construct the power spectra $P_{X}(\nu_{j})$ and $P_{Y}(\nu_{j})$ from the light curves of the energy channel and the reference band, which we denote as $x(t)$ and $y(t)$, respectively. It is important to note that the power spectra discussed here are not noise-subtracted. Using these two light curves, we also compute the cross-spectrum $C_{XY}(\nu_{j})$. To generate the power and cross-spectra, we utilize the Python-based package Stingray. All power spectra and cross-spectra are normalized using Miyamoto normalization, which is based on the fractional RMS (Miyamoto et al., 1991). After deriving power spectra and the cross-spectrum, the value of the covariance spectrum in each energy channel of interest is given by:
$Cv\left(\nu_{j}\right)=\langle x\rangle\sqrt{\frac{\Delta\nu\left(\left|\bar{C}_{XY}\left(\nu_{j}\right)\right|^{2}-n^{2}\right)}{\bar{P}_{Y}\left(\nu_{j}\right)-P_{Y,\text{ noise }}}},$ (6)
where the first term $\langle x\rangle$ is the average count in the energy channel of interest and the term $\Delta\nu_{j}$ indicates the width of this frequency range. The second term is dimensionless; we refer to it as fractional covariance.
Note that the covariance spectrum is closely related to RMS. The equation can be expressed as follows:
$Cv\left(\nu_{j}\right)=\langle x\rangle\sqrt{\gamma^{2}\left(\nu_{j}\right)}\sqrt{\left(\bar{P}_{X}\left(\nu_{j}\right)-P_{X,\text{noise}}\right)\Delta\nu_{j}},$ (7)
where $\gamma^{2}$ represents the coherence between two light curves within the specified frequency range. Since the power spectrum utilizes fractional RMS-squared normalization, the term $\sigma_{\rm X}(\nu_{j})\equiv\sqrt{\left(\bar{P}_{X}\left(\nu_{j}\right)-P_{X,\text{noise}}\right)\Delta\nu_{j}}$ indicates the fractional RMS of the energy channel within that frequency range. It is also essential to apply the background correction to this term, as given by the formula $RMS=\sqrt{P}(S+B)/S$ (Bu et al., 2015). Additionally, the error on the covariance could be derived via error propagation:
$\Delta Cv\left(\nu_{j}\right)=\sqrt{\left(\sqrt{\gamma^{2}(\nu_{j})}\cdot\sigma_{\rm X}(\nu_{j})\cdot\Delta{\rm\langle x\rangle}\right)^{2}+\left(\dfrac{\sigma_{\rm X}(\nu_{j})\cdot\langle x\rangle\cdot\Delta\gamma^{2}(\nu_{j})}{2\sqrt{\gamma^{2}(\nu_{j})}}\right)^{2}+\left(\sqrt{\gamma^{2}(\nu_{j})}\cdot\langle x\rangle\cdot\Delta\sigma_{\rm X}(\nu_{j})\right)^{2}},$ (8)
where $\Delta\sigma_{\rm X}(\nu_{j})$ is the error on the fractional rms-spectrum.